# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ Appendix B Adjacency Notions and DP Guarantees for LLM Inference
$D_{\alpha}\big{(}\mathcal{A}({\bm{q}},{\bm{R}})\|\mathcal{A}({\bm{q}},{\bm{R}}^{\prime})\big{)}\leq\rho\,\alpha\quad\text{for all}\,\,\alpha>1\,\,\text{ and all adjacent }{\bm{R}},{\bm{R}}^{\prime}\,,$ (5)
@@
Conversion formulae between various notions of indistinguishability are known [e.g. 81, Fig. 2]. For example, a $\rho$-zCDP guarantee implies $(\varepsilon,\delta)$-DP with [62, Thm. 21]
@@ B.1 Adjacency Notions in LLM Inference
This notion of adjacency is semantically meaningful and maps to intuitive expectations of a privacy guarantee. Unfortunately, it can be technically challenging to develop rigorous algorithms in this setting. In particular, we cannot assume the sensitive dataset size $|{\bm{R}}|$ to be fixed, as adjacent ${\bm{R}},{\bm{R}}^{\prime}$ are of different sizes. This makes the dataset size a privacy-sensitive quantity that needs to be protected with privacy mechanisms; see [82, §E], [14, §2.1.1], and [83].
For example, the sensitivity of the mean operation $(s_{1},\ldots,s_{n})\mapsto(1/n)\sum_{i=1}^{n}s_{i}$ is a function of the dataset size $n$, meaning that we cannot directly use it in the add-or-remove paradigm. A common workaround (used in private learning with DP-SGD) is to privatize the sum query $(s_{1},\ldots,s_{n})\mapsto\sum_{i=1}^{n}s_{i}$ and post-process it to obtain an estimator of the mean. Unfortunately, this is not possible in the setting of §3: the logit vector $\bar{\bm{\phi}}$ in Eq. 2 is actually obtained by averaging, and we cannot use the above trick here.
A common workaround to avoid these technical challenges with the add-or-remove adjacency is to use the so-called replace-by-null adjacency, where the removal of an element is simulated by swapping it with a special (problem-dependent) null element [e.g. 84, Sec. 6.2]. We instantiate the replace-by-null adjacency in the context of LLM inference in this work by choosing the null element as the empty string:
@@
Let $\perp$ denote a special null element such that the logits $\phi(y\,|\,{\bm{q}},{\bm{r}}=\perp,{\bm{x}}_{<t})=0$ are always equal to zero for any token $y\in V$, query ${\bm{q}}\in V^{*}$, and prefix ${\bm{x}}_{<t}$.999 To be fully formal, the space of all possible reference sequences ${\bm{r}}$ should be enlarged to $V^{*}\cup\{\perp\}$ to accommodate the null element $\perp$. Two sets ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{n}\}$ and ${\bm{R}}^{\prime}=\{{\bm{r}}_{1}^{\prime},\ldots,{\bm{r}}_{n}^{\prime}\}$ of equal size $n$ are said to be adjacent in the zero-out notion of adjacency if there exists an index $j\in[n]$ such that ${\bm{r}}_{i}={\bm{r}}_{i}^{\prime}$ for all $i\neq j$ and one of ${\bm{r}}_{j},{\bm{r}}_{j}^{\prime}$ equal this null element $\perp$.
Note that a zero logit vector implies next-token probabilities that are uniform over the vocabulary. As argued in §4, this is rather unrealistic as it could produce incomplete word pieces, while ${\bm{r}}=\emptyset$ as the empty string still gives reasonable next-token predictions using the pretrained model’s language and world knowledge. Further, the empty string of Definition 5 is a semantically more meaningful null element than an artificially created $\perp$ element from Definition 6. Thus, we primarily adopt the replace-by-null adjacency in this work.
@@
The replace-one (or swap) notion of adjacency is also commonly studied. Here, one element of ${\bm{R}}$ is replaced by any other arbitrary element to get ${\bm{R}}^{\prime}$. It provides an 2$\approx$ 2$\times$ 2 × stronger privacy guarantee at the same $\varepsilon$ (usually at the cost of a larger noise multiplier); see [14].
@@ B.2 Sensitivity for the Exponential Mechanism
The exponential mechanism with a target privacy level of $\rho$-zCDP selects an item $i\in\{1,\ldots,d\}$ with respective scores $[f({\bm{R}})]_{i}$ with probability [63]
@@
Recall that we wish to generate text in response to a query ${\bm{q}}$ using sensitive references ${\bm{R}}\subset V^{*}$. In this context, Definition 8 is instantiated with $f({\bm{R}})=\phi(\,\cdot\,|\,{\bm{q}},{\bm{R}},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$, where we index an element in $\mathbb{R}^{|V|}$ with a corresponding token $y\in V$.
@@ B.3 Technical Details of DP in Baselines
Amin et al. [12] claims to operate in the add-remove notion of adjacency (Definition 4), yet this is technically incorrect as we describe next. Their analysis would be correct in zero-out adjacency (Definition 6).
Error in the Analysis of [12]. Amin et al. [12, Lemma 3] claims the following (in our notation): If ${\bm{R}},{\bm{R}}^{\prime}$ are adjacent in the add-or-remove. Then, the sensitivity of the operation
@@
  • Suppose ${\bm{R}}={\bm{R}}^{\prime}\cup\{{\bm{r}}\}$, as required by add-or-remove adjacency. They claim that the sensitivity (Definition 8) of the function $f$ from Eq. 7 is
@@
Fix: Zero-Out or Replace-by-Null Adjacency. Both these issues are fixed by adopting zero-out (Definition 6) or replace-by-null (Definition 5) notions of adjacency instead of the add-or-remove adjacency. Indeed, we take $|{\bm{R}}|=|{\bm{R}}^{\prime}|$, so the first issue is fixed. Second, the dataset size can be assumed to be fixed and known in this setting (see also [14, §2.1.1]), so the sensitivity can freely depend on $|{\bm{R}}|=|{\bm{R}}^{\prime}|$.
@@
Let the dataset set $B=|{\bm{R}}|$ be fixed. Then, the sensitivity of the function $f$ defined in Eq. 7 is:
$\mathsf{sens}(f)=\begin{cases}\frac{C}{B}\,,&\text{ under zero-out adjacency}\\ \frac{2C}{B}\,,&\text{ under replace-by-null adjacency}\,.\end{cases}$
@@
Convention in our Experiments. While we use the more natural replace-by-null adjacency in our experiments, we (incorrectly) use the sensitivity bound obtained for the zero-out adjacency from Property 9. This gives Amin et al. [12] an advantage of a factor of 2 in the sensitivity. Yet, we find that InvisibleInk far outperforms their method in privacy-utility-compute tradeoffs.
@@ B.3.2 AdaPMixED: Adjacency and Data-Dependent DP Guarantees
Strictly speaking, the privacy guarantees across different adjacency notions cannot be directly compared. However, we cautiously make this comparison between the add-or-remove adjacency of the AdaPMixED and the replace-by-null adjacency of our method. This is because the semantics of these two notions of adjacency are expected to be similar [84, 14]. We can modify AdaPMixED to use the replace-by-null adjacency. We compared InvisibleInk to (AdaPMixED, add-or-remove) and (AdaPMixED, replace-by-null) and found no difference in the qualitative trends; see §E.4 for details. Hence, we stick to the original notion add-or-remove adjacency used in [18] in the main paper.
Furthermore, AdaPMixED gives data-dependent privacy guarantees. For example, the zCDP notion (Definition 3) can be phrased in the data-dependent setting under the add-or-remove adjacency as follows. An algorithm $\mathcal{A}({\bm{q}},{\bm{R}})$ satisfies $\rho({\bm{R}})$-zCDP if
@@
for all $\alpha>1$ and all datasets ${\bm{R}}^{\prime}$ that can be obtained from the given dataset ${\bm{R}}$ by removing any one of its elements; here
@@
is symmetrized Rényi divergence. This guarantee holds only for the given sensitive dataset ${\bm{R}}$. In contrast, Definition 3 must hold for all ${\bm{R}}\simeq{\bm{R}}^{\prime}$.
Thus, our data-independent guarantees are strictly more general than the data-dependent guarantees given by AdaPMixED. The latter’s guarantees are, in general, privacy-sensitive and need to be sanitized before release. In contrast, our privacy guarantees are worst-case and hold independently of the dataset.
@@ B.3.3 Adjacency Notion of AugPE
AugPE assumes the add-remove notion of adjacency (Definition 4). Their key technical step is to construct a nearest neighbor histogram. Given a set $S=({\bm{x}}_{1},\ldots,{\bm{x}}_{m})$ of candidate synthetic generations and a set $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$ of real data points, the idea is to compute the nearest neighbor histogram
@@
is the nearest neighbor of ${\bm{r}}_{j}$ from the set $S$ according to a given distance metric $\mathsf{dist}(\cdot\,,\,\cdot)$ (with ties resolved arbitrarily) and $\mathbb{I}(\cdot)$ is the indicator function which takes the value 1 when its argument is true and zero otherwise. Xie et al. [36] take $\mathsf{dist}$ as the Euclidean distance in some embedding space. This histogram $H(\cdot)\in\mathbb{R}^{m}$ is then privatized using the Gaussian mechanism. Its $\ell_{2}$-sensitivity is as follows (with an elementary proof provided for completeness).
@@
  • $1$ under the add-or-remove adjacency;
@@
Convention in our Experiments. Similar to the case of Amin et al. [12], we give AugPE an advantage by incorrectly using its best sensitivity from Property 10. In particular, we use the more natural replace-by-null adjacency in our experiments but we (incorrectly) use the AugPE’s sensitivity bound obtained for the add-or-remove adjacency from Property 10. This gives AugPE an advantage of a factor of $\sqrt{2}$ in the sensitivity. Yet, we find that InvisibleInk far outperforms AugPE in privacy-utility-compute tradeoffs.
@@ Appendix C Full Proofs and Additional Details of InvisibleInk
Below, we give the proofs of Property 1 and Theorem 2. We also give the sensitivity of DClip under different notions of adjacency below.
Proof of Property 1.
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big{(}{\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big{)}-\mathsf{clip}_{C}\big{(}{\bm{\phi}}_{1}^{\prime}-{\bm{\phi}}_{\mathsf{pub}}\big{)}\right\rVert_{\infty}$
=$\stackrel{{\scriptstyle(b)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big{(}{\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big{)}\right\rVert_{\infty}$
@@
where the last line follows because $\ell_{\infty}$ norm of any logits clipped using $\mathsf{clip}_{C}(\cdot)$ is upper bounded by $C$. ∎
Proof of Theorem 2.
Algorithm 1 consists of the application of the map $f_{\mathsf{DClip}}$, defined previously in Eq. 9, followed by the application of the exponential mechanism with a temperature of $\tau$ (in the form of softmax sampling). Note that the Top-$k+$ selection step in Algorithm 1 is agnostic to the values of the private logits ${\bm{\phi}}_{1},\dots,{\bm{\phi}}_{B}$ and therefore does not incur any additional privacy cost.
Consider the generation of one token $x_{t}$ using Algorithm 1 given arbitrary ${\bm{x}}_{<t}$. Using the definition of the exponential mechanism in §B.2 (cf. (6)), the $\rho$-zCDP guarantees for generating any one token $x_{t}$ is given by
@@
where $\mathsf{sens}(f_{\mathsf{DClip}})$ is obtained from Property 1.
To generate sequences which may contain at most $T$ tokens $x_{1},\ldots,x_{T}$, the zCDP guarantees obtained for single-token generation in Eq. 10 may be adaptively composed sequentially over all $T$ tokens. Then, the entire sequence ${\bm{x}}=(x_{1},\ldots,x_{T^{\prime}})$ with $T^{\prime}\leq T$ satisfies $\rho_{\mathsf{seq}}$-zCDP where (cf. Lemma 13(a))
@@
This completes the proof of Theorem 2. ∎
@@
Consider partitioning a dataset $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$ into $n$ disjoint batches ${\bm{R}}_{1},\ldots,{\bm{R}}_{n}$ of size $B=N/n$ (assumed an integer), i.e., ${\bm{R}}_{i}\cap{\bm{R}}_{j}=\emptyset\,\forall\,i\neq j$. Suppose further that this partition is data-independent.101010 That is, a partitioning is performed using a process that is agnostic to the content of the data. This ensures that the partition is not influenced by any patterns, information, or outcomes within the dataset. This precludes partitioning based on clustering, for instance. Suppose we generate texts ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ using Algorithm 1 such that ${\bm{x}}_{i}$ only sees sensitive examples ${\bm{R}}_{i}$ and satisfies $\rho_{\mathsf{seq}}$-zCDP individually. Then, the map $\mathcal{D}\mapsto({\bm{x}}_{1},\ldots,{\bm{x}}_{n})$ also satisfies $\rho_{\mathsf{seq}}$-zCDP.
@@
This follows from Theorem 2, and from the parallel composition for $\rho$-zCDP guarantees in Lemma 13. ∎
Note that the privacy guarantee in Corollary 11 does not depend on the number of synthetic texts generated. The number of possible generations using InvisibleInk for a fixed dataset size $N$ is thus limited only by the number of possible partitions $n\leq\lfloor N/B\rfloor$.

### v2

@@ Appendix B Adjacency Notions and DP Guarantees for LLM Inference
[…]
@@
Conversion formulae between various notions of indistinguishability are known [e.g. 92, Fig. 2]. For example, a $\rho$-zCDP guarantee implies $(\varepsilon,\delta)$-DP with [69, Thm. 21]
@@ B.1 Adjacency Notions in LLM Inference
This notion of adjacency is semantically meaningful and maps to intuitive expectations of a privacy guarantee. Unfortunately, it can be technically challenging to develop rigorous algorithms in this setting. In particular, we cannot assume the sensitive dataset size $|{\bm{R}}|$ to be fixed, as adjacent ${\bm{R}},{\bm{R}}^{\prime}$ are of different sizes. This makes the dataset size a privacy-sensitive quantity that needs to be protected with privacy mechanisms; see [93, §E], [15, §2.1.1], and [94] for further discussions.
For example, the sensitivity of the mean operation $(s_{1},\ldots,s_{n})\mapsto(1/n)\sum_{i=1}^{n}s_{i}$ is a function of the dataset size $n$, meaning that we cannot directly use it in the add-or-remove paradigm. A common workaround (used, for example, in private learning with DP-SGD) is to privatize the sum query $(s_{1},\ldots,s_{n})\mapsto\sum_{i=1}^{n}s_{i}$ and post-process it to obtain an estimator of the mean. Unfortunately, this is not possible in the setting of §3: the logit vector $\bar{\bm{\phi}}$ in Eq.˜2 is actually obtained by averaging, and we cannot use the above trick here.
A common workaround to avoid these technical challenges with the add-or-remove adjacency is to use the so-called replace-by-null adjacency, where the removal of an element is simulated by swapping it with a special (problem-dependent) null element [e.g. 70, Sec. 6.2]. We instantiate the replace-by-null adjacency in the context of LLM inference in this work by choosing the null element as the empty string:
@@
Let $\perp$ denote a special null element such that the logits $\phi(y\,|\,{\bm{q}},{\bm{r}}=\perp,{\bm{x}}_{<t})=0$ are always equal to zero for any token $y\in V$, query ${\bm{q}}\in V^{*}$, and prefix ${\bm{x}}_{<t}$. [^8] To be fully formal, the space of all possible reference sequences $\bm{r}$} should be enlarged to $V^{*}\cup\{\perp\}$ to accommodate the null element $\perp$. Two sets ${\bm{R}}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{n}\}$ and ${\bm{R}}^{\prime}=\{{\bm{r}}_{1}^{\prime},\ldots,{\bm{r}}_{n}^{\prime}\}$ of equal size $n$ are said to be adjacent in the zero-out notion of adjacency if there exists an index $j\in[n]$ such that ${\bm{r}}_{i}={\bm{r}}_{i}^{\prime}$ for all $i\neq j$ and one of ${\bm{r}}_{j},{\bm{r}}_{j}^{\prime}$ equal this null element $\perp$.
Note that a zero logit vector implies next-token probabilities that are uniform over the vocabulary. As argued in §4, this is rather unrealistic as it could produce incomplete word pieces, while $\bm{r}$$\emptyset$ as the empty string still gives reasonable next-token predictions using the pretrained model’s language and world knowledge. Further, the empty string of Definition˜5 is a semantically more meaningful null element than an artificially created $\perp$ element from Definition˜6. Thus, we primarily adopt the replace-by-null adjacency in this work.
@@
The replace-one (or swap) notion of adjacency is also commonly studied. Here, one element of $\bm{R}$} is replaced by any other arbitrary element to get $\bm{R}$$^{\prime}$. It provides an $\approx 2\times$ stronger privacy guarantee at the same $\varepsilon$ (usually at the cost of a larger noise multiplier); see [15] for a discussion.
@@ B.2 Sensitivity for the Exponential Mechanism
The exponential mechanism with a target privacy level of $\rho$-zCDP selects an item $i\in\{1,\ldots,d\}$ with respective scores $f({\bm{R}})]_{i}$ with probability [71]
@@
Recall that we wish to generate text in response to a query $\bm{q}$} using sensitive references ${\bm{R}}\subset V^{*}$. In this context, Definition˜8 is instantiated with $f({\bm{R}})=\phi(\,\cdot\,|\,{\bm{q}},{\bm{R}},{\bm{x}}_{<t})\in\mathbb{R}^{|V|}$, where we index an element in $\mathbb{R}^{|V|}$ with a corresponding token $y\in V$.
@@ B.3 Technical Details of DP in Baselines
Amin et al. [13] claims to operate in the add-remove notion of adjacency (Definition˜4), yet this is technically incorrect as we describe next. Their analysis would be correct in zero-out adjacency (Definition˜6).
Error in the Analysis of [13]. Amin et al. [13, Lemma 3] claims the following (in our notation): If ${\bm{R}},{\bm{R}}^{\prime}$ are adjacent in the add-or-remove sense, the sensitivity of the operation
@@
  • Suppose ${\bm{R}}={\bm{R}}^{\prime}\cup\{{\bm{r}}\}$, as required by add-or-remove adjacency. They claim that the sensitivity (Definition˜8) of the function $f$ from Eq.˜8 is
@@
Fix: Zero-Out or Replace-by-Null Adjacency. Both these issues are fixed by adopting zero-out (Definition˜6) or replace-by-null (Definition˜5) notions of adjacency instead of the add-or-remove adjacency. Indeed, we take $|{\bm{R}}|=|{\bm{R}}^{\prime}|$, so the first issue is fixed. Second, the dataset size can be assumed to be fixed and known in this setting (see also [15, §2.1.1]), so the sensitivity can freely depend on $|{\bm{R}}|=|{\bm{R}}^{\prime}|$.
@@
Let the dataset set $B=|{\bm{R}}|$ be fixed. Then, the sensitivity of the function $f$ defined in Eq.˜8 is:
$\mathsf{sens}(f)=\begin{cases}\frac{C}{B}\,,&\text{ under zero-out adjacency}\,,\\ \frac{2C}{B}\,,&\text{ under replace-by-null adjacency}\,.\end{cases}$
@@
Convention in our Experiments. While we use the more natural replace-by-null adjacency in our experiments, we (incorrectly) use the sensitivity bound obtained for the zero-out adjacency from Property˜9. This gives Amin et al. [13] an advantage of a factor of 2 in the sensitivity. Yet, we find that InvisibleInk far outperforms their method in privacy-utility-compute tradeoffs.
@@ B.3.2 AdaPMixED: Adjacency and Data-Dependent DP Guarantees
Strictly speaking, the privacy guarantees across different adjacency notions cannot be directly compared. Thus, we modify AdaPMixED to use the replace-by-null adjacency and report these results in the main paper.
We can also cautiously compare the add-or-remove adjacency of AdaPMixED and the replace-by-null adjacency of InvisibleInk. This is because the semantics of these two notions of adjacency are expected to be similar [70, 15]. Comparing (AdaPMixED, add-or-remove) and (AdaPMixED, replace-by-null) reveals no significant difference in the qualitative trends; see §E.5 for details.
Furthermore, AdaPMixED gives data-dependent privacy guarantees. For example, the zCDP notion (Definition˜3) can be phrased in the data-dependent setting under the replace-by-null adjacency as follows. An algorithm $\mathcal{A}({\bm{q}},{\bm{R}})$ satisfies $\rho({\bm{R}})$-zCDP if
@@
for all $\alpha>1$ and all datasets $\bm{R}$$^{\prime}$ that can be obtained from the given dataset $\bm{R}$} by replacing any one of its elements with the empty string $\emptyset$; here
@@
is symmetrized Rényi divergence. This guarantee holds only for the given sensitive dataset $\bm{R}$}. In contrast, Definition˜3 must hold for all ${\bm{R}}\simeq{\bm{R}}^{\prime}$.
Thus, our data-independent guarantees are strictly more general than the data-dependent guarantees given by AdaPMixED. The latter are, in general, privacy-sensitive and need to be sanitized before release. In contrast, our privacy guarantees are worst-case and hold independent of the dataset.
@@ B.3.3 Adjacency Notion of AugPE
AugPE assumes the add-remove notion of adjacency (Definition˜4). Their key technical step is to construct a nearest neighbor histogram. Given a set $S=({\bm{x}}_{1},\ldots,{\bm{x}}_{m})$ of candidate synthetic generations and a set $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$ of real data points, the idea is to compute the nearest neighbor histogram
@@
[…]
@@
  • 11 under the add-or-remove adjacency;
@@
Convention in our Experiments. Similar to the case of Amin et al. [13], we give AugPE an advantage by incorrectly using its best sensitivity from Property˜10. In particular, we use the more natural replace-by-null adjacency in our experiments but we (incorrectly) use the AugPE’s sensitivity bound obtained for the add-or-remove adjacency from Property˜10. This gives AugPE an advantage of a factor of $\sqrt{2}$ in the sensitivity. Yet, we find that InvisibleInk far outperforms AugPE in privacy-utility-compute tradeoffs.
@@ Appendix C Full Proofs and Additional Details of InvisibleInk
Below, we give the proofs of Property˜1 and Theorem˜2. We also give the sensitivity of DClip under different notions of adjacency below.
Proof of Property˜1.
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big({\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big)-\mathsf{clip}_{C}\big({\bm{\phi}}_{1}^{\prime}-{\bm{\phi}}_{\mathsf{pub}}\big)\right\rVert_{\infty}$
$\stackrel{{\scriptstyle(b)}}{{=}}\frac{1}{B}\left\lVert\mathsf{clip}_{C}\big({\bm{\phi}}_{1}-{\bm{\phi}}_{\mathsf{pub}}\big)\right\rVert_{\infty}$
@@
where the last line follows because $\ell_{\infty}$ norm of logits clipped using $\mathsf{clip}_{C}(\cdot)$ is bounded by $C$. ∎
Proof of Theorem˜2.
Algorithm˜1 consists of the application of the map $f_{\mathsf{DClip}}$, defined previously in Eq.˜10, followed by the application of the exponential mechanism with a temperature of $\tau$ (in the form of softmax sampling). Note that the Top-$k+$ selection step in Algorithm˜1 is agnostic to the values of the private logits ${\bm{\phi}}_{1},\dots,{\bm{\phi}}_{B}$ and therefore does not incur any additional privacy cost.
Consider the generation of one token $x_{t}$ using Algorithm˜1 given arbitrary ${\bm{x}}_{<t}$. Using the definition of the exponential mechanism in §B.2 (cf. (7)), the $\rho$-zCDP guarantees for generating any one token $x_{t}$ is given by
@@
where $\mathsf{sens}(f_{\mathsf{DClip}})$ is obtained from Property˜1.
To generate sequences which may contain at most $T$ tokens $x_{1},\ldots,x_{T}$, the zCDP guarantees obtained for single-token generation in Eq.˜11 may be adaptively composed sequentially over all $T$ tokens. Then, the entire sequence ${\bm{x}}=(x_{1},\ldots,x_{T^{\prime}})$ with $T^{\prime}\leq T$ satisfies $\rho_{\mathsf{seq}}$-zCDP where (cf. Lemma˜13(a))
@@
This completes the proof of Theorem˜2. ∎
@@
Consider partitioning a dataset $\mathcal{D}=\{{\bm{r}}_{1},\ldots,{\bm{r}}_{N}\}$ into $n$ disjoint batches ${\bm{R}}_{1},\ldots,{\bm{R}}_{n}$ of size $B=N/n$ (assumed an integer), i.e., ${\bm{R}}_{i}\cap{\bm{R}}_{j}=\emptyset\,\forall\,i\neq j$. Suppose further that this partition is data-independent. [^9] That is, a partitioning is performed using a process that is agnostic to the content of the data. This ensures that the partition is not influenced by any patterns, information, or outcomes within the dataset. This precludes partitioning based on clustering, for instance. Suppose we generate texts ${\bm{x}}_{1},\ldots,{\bm{x}}_{n}$ using Algorithm˜1 such that ${\bm{x}}_{i}$ only sees sensitive examples ${\bm{R}}_{i}$ and satisfies $\rho_{\mathsf{seq}}$-zCDP individually. Then, the map $\mathcal{D}\mapsto({\bm{x}}_{1},\ldots,{\bm{x}}_{n})$ also satisfies $\rho_{\mathsf{seq}}$-zCDP.
@@
This follows from Theorem˜2, and from the parallel composition for $\rho$-zCDP guarantees in Lemma˜13. ∎
Note that the privacy guarantee in Corollary˜11 does not depend on the number of synthetic texts generated. The number of possible generations using InvisibleInk for a fixed dataset size $N$ is thus limited only by the number of possible partitions $n\leq\lfloor N/B\rfloor$.